A salary after a 4.5% increase, if the original salary is x dollars
Answer by Maths68(1474)

Salary = x dollars
Increase =4.5% of x
Increase =(4.5/100)x
Increase =0.045x

Salary after increase = Salary Increase + Salary
Salary after increase = 0.045x + x
Salary after increase = 1.045x

Example
========
Let Salary = 4000
4.5% of 4000 = 4.5/100(4000)
4.5% of 4000 = 4.5(40)
4.5% of 4000 = 4.5(40)
4.5% of 4000 = 180
Salary after increase = 180 + 4000
Salary after increase = 4180

Now put salary in the following formula
Salary after increase = 1.045x
Salary after increase = 1.045(4000)
Salary after increase = 4180